OpenStudy (anonymous):

Factor completely: x^2 − 10x + 25

jimthompson5910 (jim_thompson5910):

Find two numbers that multiply to 25 and add to -10.

jimthompson5910 (jim_thompson5910):

What are those two numbers?

OpenStudy (anonymous):

we can use the formula a^-2ab+b^2=(a-b)^2

OpenStudy (anonymous):

FYI I JUST posted a long explanation of that formula in this thread: http://openstudy.com/study#/updates/4fe8e018e4b0bbec5cf9509c

OpenStudy (candyme):



OpenStudy (anonymous):

x^2 − 10x + 25 x^2 - 2(x)(5) + 5^2 now , a=x and b=5 so, x^2 − 10x + 25 =(x-5)^2

OpenStudy (anonymous):

okayy thanks guysss